OpenStudy (anonymous):

How do I solve 1/x=5/x-1/11

OpenStudy (anonymous):

multiply first fraction by 5 so 5 / 5x = 5 / {(x-1) / 11} so numerators are equal, so are denominators. 5x = {(x-1) / 11} 55x = x - 1 54x = -1 x = - 1 / 54 which is an odd answer unless you meant to write a different equation. if you did, re-write it using equation.